# footnotetext: File: Ahh-Raj-P1-Geo._Charc.-01-01-2026.tex, printed: 2026-1-3, 16.51

### v1

@@ Abstract.
In this article, we consider the family $\mathcal{F}(\alpha)$ defined for $\alpha\in(0,3]$ by
${\rm Re}\left(1+\frac{zf^{\prime\prime}(z)}{f^{\prime}(z)}\right)>1-\frac{\alpha}{2}\quad\text{for }z\in\mathbb{D}.$
Our primary objective is to show that this family possesses significant geometric and analytic properties, including connections with Blaschke products and the Schwarzian derivative, as well as its sharp bounds. Furthermore, we prove that if $f\in\mathcal{F}(\alpha)$, then the image $f(\mathbb{D})$ is a quasidisk. We also show that if $f\in\mathcal{F}(\alpha)$, then $\|S_{f}\|=2\alpha(2-\alpha)$. Moreover, we establish the sharp estimate $\|P_{f}\|\leq 2\alpha+1$ for the pre-Schwarzian derivative of harmonic mappings $f=h+\bar{g}\in\mathcal{F}_{\mathcal{H}}(\alpha)$, where the analytic part $h$ belongs to $\mathcal{F}(\alpha)$.
@@ 1. Introduction
$\mathcal{C}(\phi):=\bigg\{f\in\mathcal{S}:a+\frac{zf^{\prime\prime}(z)}{f^{\prime}(z)}\prec\phi\;\mbox{for}\;z\in\mathbb{D}\bigg\}$
@@
$\mathcal{S}^{*}(\phi):=\bigg\{f\in\mathcal{S}:a+\frac{zf^{\prime}(z)}{f(z)}\prec\phi\;\mbox{for}\;z\in\mathbb{D}\bigg\},$
@@
(1.1) ${\rm Re}\left(1+\frac{zf^{\prime\prime}(z)}{f^{\prime}(z)}\right)>1-\frac{\alpha}{2}\quad\text{for }z\in\mathbb{D}.$
In this work, we investigate several geometric properties of the family $\mathcal{F}(\alpha)$. The article is organized as follows. In Section 2, we develop the necessary background to characterize functions in the class $\mathcal{F}(\alpha)$ in terms of finite Blaschke products (see Theorem 2.1). In Section 3, we obtain a sharp estimate of the Schwarzian derivative for functions belonging to the family $\mathcal{F}(\alpha)$ (see Theorem 3.2). In Section 4, we discuss the properties of quasidisks; specifically, we establish a pivotal lemma (Lemma 4.1) that serves as a cornerstone for the derivation of our primary results. In Section 5, we obtain results on univalent harmonic mappings whose analytic parts belong to $\mathcal{F}(\alpha)$. Finally, in Section 6, we study pre-Schwarzian derivatives for harmonic mappings with a fixed analytic part in $\mathcal{F}(\alpha)$.

### v2

# footnotetext: File: Ahh-Raj-P1-Geo._Charc.-014-01-2026.tex, printed: 2026-1-14, 5.42

@@ Abstract.
We investigate the family $\mathcal{F}(\alpha)$, where $\alpha\in(0,3]$, of locally univalent analytic functions defined by the condition:
$\operatorname{Re}\left(1+\frac{zf^{\prime\prime}(z)}{f^{\prime}(z)}\right)>1-\frac{\alpha}{2},\quad z\in\mathbb{D}.$
Our main contributions include establishing sharp bounds for the Schwarzian and pre-Schwarzian derivatives within this family and revealing their intrinsic connection with finite Blaschke products. Furthermore, we prove that the image domains $f(\mathbb{D})$ are quasidisks. Finally, we extend these results to sense-preserving harmonic mappings $f=h+\overline{g}$ whose analytic part $h$ belongs to $\mathcal{F}(\alpha)$, and we derive optimal pre-Schwarzian norm estimates for this class.
@@ 1. Introduction
$\mathcal{C}(\phi):=\bigg\{f\in\mathcal{S}:1+\frac{zf^{\prime\prime}(z)}{f^{\prime}(z)}\prec\phi(z)\;\mbox{for}\;z\in\mathbb{D}\bigg\}$
@@
$\mathcal{S}^{*}(\phi):=\bigg\{f\in\mathcal{S}:\frac{zf^{\prime}(z)}{f(z)}\prec\phi(z)\;\mbox{for}\;z\in\mathbb{D}\bigg\},$
@@
(1.1) ${\rm Re}\left(1+\frac{zh^{\prime\prime}(z)}{h^{\prime}(z)}\right)>1-\frac{\alpha}{2}\quad\text{for }z\in\mathbb{D}.$
Our primary contribution in this paper lies in establishing sharp bounds for both the Schwarzian and pre-Schwarzian derivatives within this class, while demonstrating their fundamental relationship with finite Blaschke products. We further characterize the geometric properties of this family by proving that the image domains $f(\mathbb{D})$ constitute quasidisks, a significant result that ensures the boundary of the image is a Jordan curve with specific regularity. Finally, we broaden the scope of our study by extending these analytic findings to the realm of sense-preserving harmonic mappings $f=h+\overline{g}$; by assuming the analytic part $h$ belongs to $\mathcal{F}(\alpha)$, we successfully derive sharp pre-Schwarzian norm estimates for this more general class of functions.
The article is organized as follows. In Section 2, we develop the necessary background to characterize functions in the class $\mathcal{F}(\alpha)$ in terms of finite Blaschke products (see Theorem 2.1). In Section 3, we obtain a sharp estimate of the Schwarzian derivative for functions belonging to the family $\mathcal{F}(\alpha)$ (see Theorem 3.2). In Section 4, we discuss the properties of quasidisks; specifically, we establish a pivotal lemma (Lemma 4.1) that serves as a cornerstone for the derivation of our primary results. In Section 5, we obtain results on univalent harmonic mappings whose analytic parts belong to $\mathcal{F}(\alpha)$. Finally, in Section 6, we study pre-Schwarzian derivatives for harmonic mappings with a fixed analytic part in $\mathcal{F}(\alpha)$.
@@ 5. Results on univalent harmonic mappings with analytic part from $\mathcal{F}(\alpha)$
Figure 1. In the left panel ($\alpha=1$), we observe a smooth, convex boundary, representing the most fundamental example of a quasidisk. The middle panel ($\alpha=3$) illustrates a domain that, while no longer convex, remains a quasidisk; its boundary is a Jordan curve free of cusps, satisfying the ‘bounded turning’ property we established. Conversely, the right panel depicts a non-quasidisk domain with a sharp inward cusp, where the boundary violates the Ahlfors condition. Our results prove that functions in $\mathcal{F}(\alpha)$ effectively avoid such pathological behavior.
Figure 2. This figure illustrates the geometric evolution of image domains $f(\mathbb{D})$ in the family $\mathcal{F}(\alpha)$. Panels (i) ($\alpha=1.5$) and (ii) ($\alpha=2.0$) depict highly regular geometries, with the latter representing the classical convex case. As $\alpha$ increases to (iii) $2.3$ and (iv) $2.7$, the domains lose convexity but remain univalent quasidisks, as established in Theorem 4.1. While a general quasidisk (v) may feature a non-smooth Jordan curve satisfying the Ahlfors condition, it remains distinct from the non-quasidisk domain (6), which possesses a sharp inward cusp that violates the bounded turning property. Our results prove that for $\alpha\in(0,3]$, functions in $\mathcal{F}(\alpha)$ maintain boundary regularity and avoid such pathological singularities.
[…]
@@ 7. Estimate of Bloch constant for the class $\mathcal{F}_{\mathcal{H}}(\alpha)$
8. Geometric interpretation of the image domains for different values $\alpha\in[0,3]$.